# Low-Degree Polynomials Are Good Extractors

### v1

@@ 2.1 Low-degree polynomials extract from small families
Recall that our goal in Theorem 1 is to obtain concentration bounds for $\operatorname{bias}_{\mathbf{X}}(f)=\mathbb{E}_{x\sim\mathbf{X}}[(-1)^{f(x)}]$ with $\mathbf{X}$ an arbitrary $(n,k)$-source. As mentioned above, our simple argument combines the original result of [BHL12] for $\mathbf{X}$ uniformly distributed over $\mathbb{F}_{2}^{n}$ and an application of the leftover hash lemma.
@@
$\mathbb{E}_{f}[\operatorname{bias}_{\mathbf{X}}(f)^{t}]$ $=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\operatorname{eval}_{d}(x^{(1)})+\cdots+\operatorname{eval}_{d}(x^{(1)})=0]$
$=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg p\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0].$ (1)
@@
where $g\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ is a random degree $\leq d$ polynomial. This holds since, by Equation 1,
$\mathbb{E}_{g}[\operatorname{bias}_{L(\mathbf{X})}(g)^{t}]$ $=\Pr_{y^{(1)},\dots,y^{(t)}\sim L(\mathbf{X})}[\forall q\in\mathbb{F}_{2}[y_{1},\dots,y_{m}],\deg q\leq d:q(y^{(1)})+\cdots+q(y^{(t)})=0]$
$=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg p\leq d:p(L(x^{(1)}))+\cdots+p(L(x^{(t)}))=0]$
@@
where the last inequality uses the fact that $\deg(p\circ L)\leq d$ (with $p\circ L$ denoting composition), as $L$ is linear.
We are almost done. Informally, since $L(\mathbf{X})\approx\mathbf{U}_{m}$, the uniform distribution on $\mathbb{F}_{2}^{m}$, it is easy to show that $\mathbb{E}_{g}[\operatorname{bias}_{L(\mathbf{X})}(g)^{t}]\approx\mathbb{E}_{g}[\operatorname{bias}(g)^{t}]$. Moreover, we can upper bound $\mathbb{E}_{g}[\operatorname{bias}(g)^{t}]$ appropriately via the known concentration bound from [BHL12]. Combining this with Equation 2 yields the desired upper bound on $\mathbb{E}_{f}[\operatorname{bias}_{\mathbf{X}}(f)^{t}]$, which we translate into a concentration bound on $\operatorname{bias}_{\mathbf{X}}(f)$ via Markov’s inequality.
@@ 2.2.1 Low-degree polynomials disperse from sumset sources
We discuss the approach behind Theorem 2 by focusing first on the simpler task of dispersing. In this case, we wish to show that a random degree $\leq d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ will satisfy $f(\mathsf{supp}(\mathbf{W}))=\{0,1\}$ simultaneously for all $(n,k)$-sumset sources $\mathbf{W}$, with $k=O(dn^{2/d})$. Then, we discuss the necessary modifications to obtain sumset extraction with polynomial error $n^{-\Omega(1)}$.
First, as usual, we only need to focus on sumset sources $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ where $\mathbf{X}$ and $\mathbf{Y}$ are independent flat $(n,k)$-sources. Denote the supports of $\mathbf{X}$ and $\mathbf{Y}$ by $\mathcal{X}$ and $\mathcal{Y}$, respectively, which have size at least $2^{k}$. Then, the probability that $f$ is identically 00 on $\mathcal{X}+\mathcal{Y}$ is
$\Pr_{f}[f(\mathcal{X}+\mathcal{Y})\equiv 0]\leq 2^{-\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))}.$ (3)
This is because we may write $f(x)=\langle v,\operatorname{eval}_{d}(x)\rangle$ for a uniformly random vector $v\in\mathbb{F}_{2}^{n}$, and so (i) $f(x)$ is uniformly distributed over $\mathbb{F}_{2}$ for any fixed $x$, and (ii) $f(x)$ and $f(y)$ are independent (and uniformly distributed) if $\operatorname{eval}_{d}(x)$ and $\operatorname{eval}_{d}(y)$ are linearly independent.
Given Equation 3, it is clear that we must understand $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))$. If this quantity is suitably large, then the probability that $f$ is constant on any such sumset $\mathcal{X}+\mathcal{Y}$ is small, and we could hope to survive a union bound over all choices of $\mathcal{X}$ and $\mathcal{Y}$. However, this strategy cannot directly work, because $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))$ will be at most $\binom{n}{\leq d}$ while there are $\binom{2^{n}}{2^{k}}^{2}\geq 2^{2(n-k)2^{k}}\gg 2^{\binom{n}{\leq d}}$ choices for $\mathcal{X}$ and $\mathcal{Y}$.
A possible way to overcome the barrier to the application of the union bound above is to show that there exist appropriately small subsets $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ such that $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}^{\prime}+\mathcal{Y}^{\prime}))$ is still large. If this was the case, we could then just apply the union bound over all possible choices of the now much smaller sets $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$. We can make this approach work by proving the following:
@@
We sketch how Claim 1 can be applied to obtain the desired result. Setting $A=\mathcal{X}$ and $B=\mathcal{Y}$, we obtain $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ of size about $\sqrt{\binom{k}{\leq d}}$ such that $\operatorname{eval}_{d}(\mathcal{X}^{\prime}+\mathcal{Y}^{\prime})$ has rank about $\binom{k}{\leq d}$. By Equation 3, this means that the probability that $f$ is identically 00 on $\mathcal{X}^{\prime}+\mathcal{Y}^{\prime}$, and hence on $\mathcal{X}+\mathcal{Y}$, is at most about $2^{-\binom{k}{\leq d}}$. But now, crucially, we only have to carry out a union bound over the at most about $\binom{2^{n}}{k^{d/2}}^{2}$ choices for $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$, which is approximately $2^{nk^{d/2}}\ll 2^{\binom{k}{\leq d}}$ when $k=\Omega(dn^{2/d})$. Note that this strategy would not have worked if $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$ were instead of size close to $\binom{k}{\leq d}$.
@@
Applying this same rank-preserving surjection in the setting of Claim 1 is not guaranteed to work, because now the subset that witnesses the $\binom{k}{\leq d}$ rank lower bound must be a sumset $A+B$ with $|A|,|B|\approx\sqrt{\binom{k}{\leq d}}$. We overcome this by observing that if we are fine with worse constants, then we can replace the complicated rank-preserving surjection from [KS05, BHL12] with a linear map, guaranteed by the leftover hash lemma. In fact, by the leftover hash lemma, there exists a linear map $L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{k^{\prime}}$, for $k^{\prime}=\Omega(k)$, which is surjective on both $\mathcal{X}$ and $\mathcal{Y}$. We choose $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ such that $L(\mathcal{X}^{\prime})=L(\mathcal{Y}^{\prime})=\mathcal{B}^{k^{\prime}}_{d/2}(0)$, the radius-$d/2$ Hamming ball in $\mathbb{F}_{2}^{k^{\prime}}$ centered at 00. Then, the linearity of $L$ plus basic properties of evaluation vectors allows us to conclude that
$\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}^{\prime}+\mathcal{Y}^{\prime}))\geq\operatorname{rank}(\operatorname{eval}_{d}(L(\mathcal{X}^{\prime}+\mathcal{Y}^{\prime})))=\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{k^{\prime}}_{d/2}(0)+\mathcal{B}^{k^{\prime}}_{d/2}(0)))=\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{k^{\prime}}_{d}(0))),$
@@ 2.2.2 From dispersers to extractors via random convex combinations
The argument discussed above shows that a random degree $\leq d$ polynomial is a $k$-sumset disperser for min-entropy $k=O(dn^{2/d})$ with high probability. It remains to see how we can extend this to get sumset extraction for similar min-entropy $k$ with polynomial error $\varepsilon=n^{-\Omega(1)}$.
Our first observation is that if $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ is a sumset source with flat $\mathbf{X}$ and $\mathbf{Y}$ whose supports $\mathcal{X}$ and $\mathcal{Y}$ satisfy $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))=|\mathcal{X}|\cdot|\mathcal{Y}|$, then
$\Pr[f(\mathbf{W})=0]\approx 1/2$
holds with high probability over the choice of a random degree $\leq d$ polynomial $f$. In other words, $f(\mathbf{W})$ is close (in statistical distance) to uniformly distributed over $\{0,1\}$ with high probability over the choice of $f$. This happens because, under the conditions above, $\mathbf{W}$ is a flat source and all the vectors in $\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y})$ are linearly independent, which means that the bits $(f(w)=\langle v,\operatorname{eval}_{d}(w)\rangle)_{w\in\mathcal{X}+\mathcal{Y}}$ are independent and uniformly distributed when $v$ is a uniformly random vector over $\mathbb{F}_{2}^{\binom{n}{\leq d}}$.
Set $k=\Omega(dn^{\frac{2(1+\delta)}{d}})$ for an arbitrarily small constant $\delta>0$ and call a sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ special if $\mathbf{X}$ and $\mathbf{Y}$ are flat, $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))=|\mathcal{X}|\cdot|\mathcal{Y}|$, and $|\mathcal{X}|,|\mathcal{Y}|\approx\sqrt{\binom{k}{\leq d}}$. Combining the previous paragraph with a union bound over the choices of $\mathcal{X}$ and $\mathcal{Y}$ (analogous to the one in Section 2.2.1) shows that a random degree $\leq d$ polynomial will be, with high probability, an $(\varepsilon=n^{-\Omega(1)})$-extractor for the class of special sumset sources.
Of course, most $(n,k)$-sumset sources $\mathbf{W}$ are far from being special. We overcome this by showing that every $(n,k)$-sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ with flat $\mathbf{X}$ and $\mathbf{Y}$ is $2^{-\Omega(k)}$-close to some convex combination of special sumset sources. Combining this with the observations above lets us conclude that a random degree $\leq d$ polynomial will be a $(k,\varepsilon^{\prime}=2^{-\Omega(k)}+n^{-\Omega(1)}=n^{-\Omega(1)})$-sumset extractor with high probability.
@@
We consider an alternative way of (approximately) sampling from $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ – the convex combination will be implicit in this sampling procedure. The idea is to first sample uniformly random subsets $\mathcal{X}^{\prime}\subseteq\mathcal{X}=\mathbb{F}_{2}^{k}$ and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}=\mathbb{F}_{2}^{k}$ each of size $\binom{k/3}{d/2}$ (which is very roughly $\sqrt{\binom{k}{\leq d}}$), and then sample $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ uniformly at random from $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$, respectively. If $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$ are sampled independently, then it is not hard to show that $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is distributed exactly like $\mathbf{W}$. However, we would like to claim that the resulting sumset source $\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ will be special with high probability over the choice of subsets $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$. To this end, we do not sample the subsets $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$ independently from each other, but rather couple the randomness used in their sampling carefully. This coupling will ensure that $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is always a special sumset source for any fixing of $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$, while we still have $\mathbf{W}^{\prime}\approx_{2^{-\Omega(k)}}\mathbf{W}$.
To sample the (correlated) random subsets $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$, we proceed as follows. Let $B_{1}=\{u_{1},\dots,u_{t}\}$ be the set of weight-$d/2$ vectors supported on $\{1,\dots,k/3\}$, and let $B_{2}=\{v_{1},\dots,v_{t}\}$ be the set of weight-$d/2$ vectors supported on $\{2k/3+1,\dots,k\}$. Since any two vectors $u_{i}$ and $v_{j}$ have disjoint supports and are non-zero, each sum $u_{i}+v_{j}$ is a distinct non-zero vector in the radius-$d$ Hamming ball, and so
@@
We couple the sampling of $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$ by choosing a uniformly random invertible matrix $\mathbf{L}\in\mathbb{F}_{2}^{k\times k}$ and setting $\mathcal{X}^{\prime}=\mathbf{L}B_{1}=\{\mathbf{L}u_{1},\dots,\mathbf{L}u_{t}\}$ and $\mathcal{Y}^{\prime}=\mathbf{L}B_{2}=\{\mathbf{L}v_{1},\dots,\mathbf{L}v_{t}\}$. Now, because $\mathbf{L}$ is invertible, we know that
$\operatorname{rank}\operatorname{eval}_{d}(\mathcal{X}^{\prime}+\mathcal{Y}^{\prime})=\operatorname{rank}\operatorname{eval}_{d}(B_{1}+B_{2})=|B_{1}|\cdot|B_{2}|=|\mathcal{X}^{\prime}|\cdot|\mathcal{Y}^{\prime}|.$
@@
Therefore, if $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ are sampled independently and uniformly at random from $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$, respectively, then $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is a special sumset source, as desired.

### v2

@@ 2.1 Low-degree polynomials extract from small families
Recall that our goal in Theorem 1 is to obtain concentration bounds for $|\operatorname{bias}_{\mathbf{X}}(f)|=|\mathbb{E}_{x\sim\mathbf{X}}[(-1)^{f(x)}]|$ with $\mathbf{X}$ an arbitrary $(n,k)$-source. Our simple argument combines the original result of [BHL12] for $\mathbf{X}$ uniformly distributed over $\mathbb{F}_{2}^{n}$ (Theorem ) and an application of the leftover hash lemma.
@@
$\mathbb{E}_{f}[\operatorname{bias}_{\mathbf{X}}(f)^{t}]$ $=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\operatorname{eval}_{d}(x^{(1)})+\cdots+\operatorname{eval}_{d}(x^{(t)})=0]$
$=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2},\deg p\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0].$ (1)
@@
where $g:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}$ is a random degree $d$ polynomial. This holds since, by Equation 1,
$\mathbb{E}_{g}[\operatorname{bias}_{L(\mathbf{X})}(g)^{t}]$ $=\Pr_{y^{(1)},\dots,y^{(t)}\sim L(\mathbf{X})}[\forall q:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2},\deg q\leq d:q(y^{(1)})+\cdots+q(y^{(t)})=0]$
$=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall q:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2},\deg q\leq d:q(L(x^{(1)}))+\cdots+q(L(x^{(t)}))=0]$
@@
where the inequality uses the fact that $q\circ L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ has degree at most $d$ (with $q\circ L$ denoting composition), as $L$ is linear.
We are almost done. Informally, since $L(\mathbf{X})\approx\mathbf{U}_{m}$, it is easy to show that $\mathbb{E}_{g}[\operatorname{bias}_{L(\mathbf{X})}(g)^{t}]\approx\mathbb{E}_{g}[\operatorname{bias}(g)^{t}]$. Moreover, we can upper bound $\mathbb{E}_{g}[\operatorname{bias}(g)^{t}]$ appropriately via the known concentration bound from [BHL12]. Combining this with Equation 2 yields the desired upper bound on $\mathbb{E}_{f}[\operatorname{bias}_{\mathbf{X}}(f)^{t}]$, which we translate into a concentration bound on $|\operatorname{bias}_{\mathbf{X}}(f)|$ via Markov’s inequality.
@@ 2.2 Low-degree polynomials extract from sumset sources
Next, we discuss the approach behind Theorem 2. For simplicity, we focus here on the case of even degree $d$, and note that it is trivial to extend to odd $d$ (at a slight loss in parameters).
@@ 2.2.1 Low-degree polynomials disperse from sumset sources
As a warm-up, we first consider the simpler task of dispersing. In this case, we wish to show that a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ will satisfy $f(\mathsf{supp}(\mathbf{W}))=\{0,1\}$ simultaneously for all $(n,k)$-sumset sources $\mathbf{W}$ with $k=O(dn^{2/d})$. Then, we discuss the necessary modifications to obtain sumset extraction with arbitrary error $\varepsilon>0$.
First, as usual, we only need to focus on sumset sources $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ where $\mathbf{X}$ and $\mathbf{Y}$ are independent flat $(n,k)$-sources. Denote the supports of $\mathbf{X}$ and $\mathbf{Y}$ by $X$ and $Y$, respectively, which have size $2^{k}$. Then, the probability that $f$ is identically 00 on $X+Y$ is
$\Pr_{f}[f(X+Y)\equiv 0]\leq 2^{-\operatorname{rank}(\operatorname{eval}_{d}(X+Y))}.$ (3)
This is because we may write $f(x)=\langle v,\operatorname{eval}_{d}(x)\rangle$ for a uniformly random vector $v\in\mathbb{F}_{2}^{\binom{n}{\leq d}}$, and so (i) $f(x)$ is uniformly distributed over $\mathbb{F}_{2}$ for any fixed nonzero $x$, and (ii) $f(x)$ and $f(y)$ are independent (and uniformly distributed) if $\operatorname{eval}_{d}(x)$ and $\operatorname{eval}_{d}(y)$ are linearly independent.
Given Equation 3, it is clear that we must understand $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))$. If this quantity is suitably large, then the probability that $f$ is constant on any such sumset $X+Y$ is small, and we could hope to survive a union bound over all choices of $X$ and $Y$. However, this strategy cannot directly work, because $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))$ will be at most $\binom{n}{\leq d}$ while there are $\binom{2^{n}}{2^{k}}^{2}\geq 2^{2(n-k)2^{k}}\gg 2^{\binom{n}{\leq d}}$ choices for $X$ and $Y$.
A possible way to overcome the barrier to the application of the union bound above is to show that there exist appropriately small subsets $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $\operatorname{rank}(\operatorname{eval}_{d}(X^{\prime}+Y^{\prime}))$ is still large. If this were the case, we could then just apply the union bound over all possible choices of the now much smaller sets $X^{\prime}$ and $Y^{\prime}$. We can make this approach work by proving the following:
@@
We sketch how Claim 1 can be applied to obtain the desired result. Setting $A=X$ and $B=Y$, we obtain $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ of size about $\sqrt{\binom{k}{\leq d}}$ such that $\operatorname{eval}_{d}(X^{\prime}+Y^{\prime})$ has rank about $\binom{k}{\leq d}$. By Equation 3, this means that the probability that $f$ is identically 00 on $X^{\prime}+Y^{\prime}$, and hence on $X+Y$, is at most about $2^{-\binom{k}{\leq d}}$. But now, crucially, we only have to carry out a union bound over the at most about $\binom{2^{n}}{k^{d/2}}^{2}$ choices for $X^{\prime}$ and $Y^{\prime}$, which is approximately $2^{nk^{d/2}}\ll 2^{\binom{k}{\leq d}}$ when $k\geq Cdn^{2/d}$ for some large enough constant $C>0$. Note that this strategy would not have worked if $X^{\prime}$ and $Y^{\prime}$ were instead of size close to $\binom{k}{\leq d}$.
@@
Applying this same rank-preserving surjection in the setting of Claim 1 is not guaranteed to work, because now the subset that witnesses the $\binom{k}{\leq d}$ rank lower bound must be a sumset $A+B$ with $|A|,|B|\approx\sqrt{\binom{k}{\leq d}}$. We overcome this by observing that if we are fine with worse constants, then we can replace the complicated rank-preserving surjection from [KS05, BHL12] with a linear map, guaranteed by the leftover hash lemma. In fact, by the leftover hash lemma, there exists a linear map $L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{k^{\prime}}$, for $k^{\prime}=\Omega(k)$, which is surjective on both $X$ and $Y$. We choose $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $L(X^{\prime})=L(Y^{\prime})=\mathcal{B}^{k^{\prime}}_{d/2}(0)$, the radius-$d/2$ Hamming ball in $\mathbb{F}_{2}^{k^{\prime}}$ centered at 00. Then, the linearity of $L$ plus basic properties of evaluation vectors allows us to conclude that
$\operatorname{rank}(\operatorname{eval}_{d}(X^{\prime}+Y^{\prime}))$ $\geq\operatorname{rank}(\operatorname{eval}_{d}(L(X^{\prime}+Y^{\prime})))$
$=\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{k^{\prime}}_{d/2}(0)+\mathcal{B}^{k^{\prime}}_{d/2}(0)))=\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{k^{\prime}}_{d}(0))),$
@@ 2.2.2 From dispersers to extractors via random convex combinations
The argument discussed above shows that a random degree $d$ polynomial is a $k$-sumset disperser for min-entropy $k=O(dn^{2/d})$ with high probability. It remains to see how we can extend this to get sumset extraction for similar min-entropy $k$ with arbitrary error $\varepsilon>0$.
Our first observation is that if $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ is a sumset source with flat $\mathbf{X}$ and $\mathbf{Y}$ whose supports $X$ and $Y$ satisfy $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))=|X|\cdot|Y|$, then $\Pr[f(\mathbf{W})=0]\approx 1/2$ holds with high probability over the choice of a random degree $d$ polynomial $f$. In other words, $f(\mathbf{W})$ is close (in statistical distance) to uniformly distributed over $\{0,1\}$ with high probability over the choice of $f$ [^13] The exact shape of the “high probability” comes from a standard Chernoff bound. This happens because, under the conditions above, $\mathbf{W}$ is a flat source and all the vectors in $\operatorname{eval}_{d}(X+Y)$ are linearly independent, which means that the bits $(f(w)=\langle v,\operatorname{eval}_{d}(w)\rangle)_{w\in X+Y}$ are independent and uniformly distributed when $v$ is a uniformly random vector over $\mathbb{F}_{2}^{\binom{n}{\leq d}}$.
Now, set $k=Cd(n/\varepsilon^{2})^{2/d}$ for a large enough constant $C>0$, and call a sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ special if $\mathbf{X}$ and $\mathbf{Y}$ are flat, $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))=|X|\cdot|Y|$, and $|X|,|Y|\approx\sqrt{\binom{k}{\leq d}}$. Combining the previous paragraph with a union bound over the choices of $X$ and $Y$ (analogous to the one in Section 2.2.1) shows that a random degree $d$ polynomial will be, with high probability, an $\varepsilon$-extractor for the class of special sumset sources.
Of course, most sumset sources are far from special. We overcome this by showing that every $(n,k)$-sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ with flat $\mathbf{X}$ and $\mathbf{Y}$ is $2^{-\Omega(k)}$-close to a convex combination of special sumset sources. Combining this with the observations above lets us conclude that a random degree $d$ polynomial will be a $(k,\varepsilon^{\prime}=2^{-\Omega(k)}+\varepsilon\approx\varepsilon)$-sumset extractor with high probability.
@@
We consider an alternative way of (approximately) sampling from $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ – the convex combination will be implicit in this sampling procedure. The idea is to first sample uniformly random subsets $X^{\prime}\subseteq X=\mathbb{F}_{2}^{k}$ and $Y^{\prime}\subseteq Y=\mathbb{F}_{2}^{k}$ each of size $\binom{k/3}{d/2}$ (which is very roughly $\sqrt{\binom{k}{\leq d}}$), and then sample $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ uniformly at random from $X^{\prime}$ and $Y^{\prime}$, respectively. If $X^{\prime}$ and $Y^{\prime}$ are sampled independently, then it is not hard to show that $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is distributed exactly like $\mathbf{W}$. However, we would like to claim that the resulting sumset source $\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ will be special with high probability over the choice of subsets $X^{\prime}$ and $Y^{\prime}$. To this end, we do not sample the subsets $X^{\prime}$ and $Y^{\prime}$ independently from each other, but rather couple the randomness used in their sampling carefully. This coupling will ensure that $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is always a special sumset source for any fixing of $X^{\prime}$ and $Y^{\prime}$, while we still have $\mathbf{W}^{\prime}\approx_{2^{-\Omega(k)}}\mathbf{W}$.
To sample the (correlated) random subsets $X^{\prime}$ and $Y^{\prime}$, we proceed as follows. Let $B_{1}=\{u_{1},\dots,u_{t}\}$ be the set of weight-$d/2$ vectors supported on $\{1,\dots,k/3\}$, and let $B_{2}=\{v_{1},\dots,v_{t}\}$ be the set of weight-$d/2$ vectors supported on $\{2k/3+1,\dots,k\}$. Since any two vectors $u_{i}$ and $v_{j}$ have disjoint supports and are non-zero, each sum $u_{i}+v_{j}$ is a distinct non-zero vector in the radius-$d$ Hamming ball, and so
@@
We couple the sampling of $X^{\prime}$ and $Y^{\prime}$ by choosing a uniformly random invertible matrix $\mathbf{L}\in\mathbb{F}_{2}^{k\times k}$ and setting $X^{\prime}=\mathbf{L}B_{1}=\{\mathbf{L}u_{1},\dots,\mathbf{L}u_{t}\}$ and $Y^{\prime}=\mathbf{L}B_{2}=\{\mathbf{L}v_{1},\dots,\mathbf{L}v_{t}\}$. Now, because $\mathbf{L}$ is invertible, we know that
$\operatorname{rank}\operatorname{eval}_{d}(X^{\prime}+Y^{\prime})=\operatorname{rank}\operatorname{eval}_{d}(B_{1}+B_{2})=|B_{1}|\cdot|B_{2}|=|X^{\prime}|\cdot|Y^{\prime}|.$
@@
Therefore, if $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ are sampled independently and uniformly at random from $X^{\prime}$ and $Y^{\prime}$, respectively, then $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is a special sumset source, as desired.